Use the method of substitution to solve the system of linear equations.
2x-y=6
7x-3y=23

Answer by stanbon(75887)

Use the method of substitution to solve the system of linear equations.
2x-y=6
7x-3y=23
------
y = 2x-6
Substitute for "y" and solve for "x":
7x - 3(2x-6) = 23
7x -6x +18 = 23
x = 41
----
Solve for "y":
y = 2x-6
y = 2*41-6
y = 82-6
y = 76
==================
cheers,
Stan H.
==================